OpenStudy (anonymous):

(9+1)^(1/2) Find the first 4 terms of this using binomial expansion,

OpenStudy (anonymous):

\[\sqrt{10}\]?

OpenStudy (anonymous):

It's like (1/2)C0(a)^(1/2)(b)^(0) + (1/2)C1(a)^(-1/2)(b)^(1) and so on for 4 terms. I just don't know how to combine it.

OpenStudy (anonymous):

oooh i see, maybe like \[(1+9)^{\frac{1}{2}}=1+\frac{1}{2}\times 9+\frac{\frac{1}{2}\times -\frac{1}{2}}{2!}9^2+..\]

OpenStudy (amistre64):

\[(1+a)^n=1+na+\frac{n(n-1)}{2!}a^2+\frac{n(n-1)(n-2)}{3!}a^3\] if i remember it correctly

OpenStudy (anonymous):

yes that is it i am sure (more or less) but because you are dealing with fractional exponent it goes on forever. and needs to look like \[(1+x)^n\]

OpenStudy (anonymous):

ummm. wut?

OpenStudy (amistre64):

your going to have to be more specific than that ....

OpenStudy (anonymous):

how did you get that string of numbers?

OpenStudy (amistre64):

its the definition of a binomial expansion for rational exponents

OpenStudy (amistre64):

im not quite sure how they come up with it yet, but i recall seeing it as such

OpenStudy (anonymous):

that is not going to work for fractional exponents

OpenStudy (anonymous):

read here for a brief explanation

OpenStudy (amistre64):

\[\binom{n}{1}=\frac{n}{1!}\] \[\binom{n}{2}=\frac{n(n-1)}{2!}\] \[\binom{n}{3}=\frac{n(n-1)(n-2)}{3!}\] etc...

OpenStudy (amistre64):

those are textbook definitions i beleive

OpenStudy (anonymous):

could you expand of (9+1)^(1/2) to 4 terms? I think that'd make more sense to me than just variables O_Ollll

OpenStudy (amistre64):

satellite posted it above

OpenStudy (amistre64):

4th post down